OpenStudy (anonymous):

How does 3^2 X 3^3 = 243? someone help..

OpenStudy (anonymous):

\[3^{2}\times3^{3}\]

OpenStudy (anonymous):

\[x^{a} \times x^b = x^{a+b}\]

OpenStudy (anonymous):

^

OpenStudy (anonymous):

3^2=9 right and 3^3= 27 soo whats 27 times 9=

OpenStudy (anonymous):

243 is correct